cdelk20:

Harry can rake the leaves in the yard 8 hours faster than his little brother Jimmy can. If they work together, they can complete the job in 3 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Jimmy to complete this job on his own.

tigerlover:

Hello and Welcome to QuestionCove! What do you think you should provide in this answer? :)

cdelk20:

what

Vocaloid:

@angle whenever you have a spare moment would you mind looking at this

AngeI:

1/x + 1/(x - 8) = 1/3 3(x - 8) + 3x = x(x - 8) x^2 - 14x + 24 = 0 (x - 12)(x - 2) = 0 so x=2 and x=12 If Harry is 8 times faster which one has to be Harry?

AngeI:

@cdelk20